OpenStudy (anonymous):

If the width of a rectangle is 8 inches and its perimetter is 44 inches what is the lengthe of the tirangle?

OpenStudy (anonymous):

8 x 44 = 352 inches right?

OpenStudy (anonymous):

so the 352 would be it's length right?

OpenStudy (ghazi):

no...you have given width of rectangle and asked for triangle.... lol

OpenStudy (anonymous):

oh wow.

OpenStudy (anonymous):

It says that the perimeter of a rectrangle equals the sum of the lengths its four sides. If the width is 8 inches and its perimeter is 44 inches , find the leght of the rectangle.

OpenStudy (ghazi):

2(length+8)=44 2*L+16=44 2L=44-16=28 L=14

OpenStudy (ghazi):

2*(length+breadth)= perimeter

OpenStudy (anonymous):

Oh I see now, Thank you

OpenStudy (ghazi):

:)